Circular No. 8632

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2005 W2 (CHRISTENSEN) A comet with a slightly asymmetric 20′′ coma elongated toward p.a.

230^{o}–260^{o} has been found by E. J. Christensen on Catalina Sky Survey images taken with the 0.68-m Schmidt telescope (discovery observation below). Others making follow-up observations that noted the cometary appearance include P. Birtwhistle (Great Shefford, Berkshire, U.K., 0.4-m reflector, Nov. 20.8 UT; 12′′ coma with a brighter 3′′ center and a possible faint extension ∼ 14′′ long in p.a. 55^{o}), J. McGaha (Tucson, AZ, 0.62-m reflector, Nov. 21.2; uniform, round 10′′ coma, no elongation), and J. Young (Table Mtn., 0.61-m reflector, Nov. 21.3, thin cirrus and moonlight; bright, diffuse 8′′ coma with faint outer area extending 16′′ to the northeast, and a second poorly defined extension seen around p.a. 240^{o} in stacked images). Further astrometry is given on MPEC 2005-W24, together with parabolic orbital elements (T = 2005 June 1.107 TT, q = 3.17566 AU, ω = 33^{o}.786, Ω = 348^{o}.766, i = 11^{o}.836, equinox 2000.0), and an ephemeris.

2005 UT α_{2000} δ_{2000} Mag. Observer
Nov. 20.27730 3^{h}47^{m}21.^{s}19 +35^{o}52′46′′.3 17.1 Christensen

SUPERNOVAE 2005kg, 2005kh, 2005ki, 2005kj, 2005kk Five apparent supernovae have been reported from unfiltered CCD im-

ages: 2005kj by the “Nearby Supernova Factory” collaboration via Palomar NEAT frames (cf. CBET 296), and the rest by E. Lee, N. J. Ponticello, and W. Li via LOSS/KAIT (cf. IAUC 8630).

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005kg Nov. 17.09 21^{h}11^{m}39.^{s}15 −23^{o}10′20′′.0 17.0 1′′.6 W, 2′′.7 S
2005kh Nov. 17.54 10 01 25.41 +15 46 49.4 18.7 7′′.6 W, 37′′ N
2005ki Nov. 18.57 10 40 28.22 + 9 12 08.4 18.0 2′′.0 W, 71′′ N
2005kj Nov. 17.5 8 40 09.18 − 5 36 02.2 16.3  —
2005kk Nov. 17.50 10 39 40.02 +25 19 25.1 17.7 13′′.0 E, 4′′.0 N

Additional approximate KAIT magnitudes: 2005kg in MCG –04-50-7, Oct. 22.12 UT, [19.0; Nov. 18.08, 17.2. SN 2005kh in NGC 3094, May 13.20, [19.5; Nov. 18.50, 18.6. SN 2005ki in NGC 3332, May 13.20, [19.5; Nov. 19.58, 17.7. 2005kk in NGC 3323, Apr. 6.18, [19.5; Nov. 21.52, 17.7. SNe 2005kg and 2005ki are type-Ia supernovae, while 2005kh is of type II and 2005kj is of type IIn (cf. CBETs 295, 296, 299).

2005 November 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green